OpenStudy (anonymous):

PLEASE HELP! D: 2. What are the zeroes of the function? Graph the function y = (x + 3)(x – 4)(x – 5). 3, –4, 5 –3, 4, 5 3, –4, –5 –3, 4, –5

hero (hero):

The zeroes of the function occur where x + 3 = 0 x - 4 = 0 x - 5 = 0 Solve for x in each

OpenStudy (kuoministers):

Think... What makes x zero? ill do x + 3 for you to make X + 3 zero we need minus 3 yes? So the answer for that is -3

hero (hero):

Bro, don't just give it away. What I said was enough.

OpenStudy (anonymous):

Thank you both of you! :) :)